OpenStudy (anonymous):

Find the equation of the tangent line at x = 2 for f(x) = 4 + x - 2x^2 - 3x^3. Write the answer in the form of y=mx+b.

OpenStudy (campbell_st):

do you know the 1st derivative.... f'(x)...?

OpenStudy (anonymous):

is it 2*2x^2-1= 4x

OpenStudy (campbell_st):

not the derivative f'(x) = 1 -4x - 9x^2 this is the equation of the slope of the tangent at any point on the curve you have the specific case of x = 2 so evaluate f'(2) which means substitute x = 2 into the derivative and get a number answer. you will also need to point on the curve... so substitute x =2 into the original equation to find y. then you will have, m x and y substitute into y = mx + b and solve for b... this will be the equation of the tangent. hope this helps

OpenStudy (anonymous):

im getting the mx but when I solve for b, it is not coming out right. im getting 17

OpenStudy (campbell_st):

ok, what did you get for f'(2) ...?